OpenStudy (anonymous):

integral of 1/((x+5)(sqrt(x^2+5x)))

OpenStudy (kkutie7):

\[\int\limits_{}^{}\frac{1}{((x+5)(\sqrt{x^{2}+5}))}\] here this is a bit easier to see.

OpenStudy (kkutie7):

\[\int\limits_{}^{}\frac{(x^{2}+5)^{-1/2}}{x+5}\] another way to write it.

OpenStudy (kkutie7):

substitution? maybe? \[u=x^{2}+5\] \[du=2xdx\] \[\frac{du}{2x}=dx\]

OpenStudy (kkutie7):

\[\int\limits_{}^{}\frac{(u)^{-1/2}}{x+5}\]

OpenStudy (kkutie7):

I have to get going or else I'd stay and try to help. I would recommend this http://www.wolframalpha.com/input/?i=intigral+of+1%2F%28%28x%2B5%29%28sqrt%28x^2%2B5x%29%29%29 for you to check your work.

OpenStudy (caozeyuan):

@Kkutie7, it is x^2+5x, you are missing a x

OpenStudy (dumbcow):

try this substitution \[x+5 = 5 \sec^2 u\] \[dx = 10 \sec^2 u \tan u du\] \[\rightarrow \int\limits \frac{10 \sec^2 u \tan u}{(5 \sec^2 u) \sqrt{5 \sec^2 u} \sqrt{5\sec^2 u - 5}} du\] \[ = \frac{2}{5} \int\limits \cos u du\] \[= \frac{2}{5} \sin u +C\] From substitution we know that \[\cos^2 u = \frac{5}{x+5}\] \[\rightarrow \sin^2 u = 1- \frac{5}{x+5} = \frac{x}{x+5}\] final answer: \[\rightarrow \frac{2}{5} \sqrt{\frac{x}{x+5}} +C\]

OpenStudy (caozeyuan):

best response of the day!@dumbcow

OpenStudy (dumbcow):

sorry for the complex substitution, i was trying to save steps you can get same result by doing w^2 = x+5 2w dw = dx then do w = sqrt5 sec u dw = sqrt5 sec u tan u

OpenStudy (dumbcow):

@caozeyuan , well thank you :)